OpenStudy (anonymous):

Write an equation in point-slope form that passes through the points (5,2) (-3,11)

OpenStudy (anonymous):

Point slope from is (y-y1)=m(x-x1) where m is the slope and (x1,y1) is a point of the line..... So find m, the slope, by using the formula for slope \[\frac{ y_2-y_1}{ x_2-x_1 }\] So in this case the slope is -9/8.... Now you can use point-slope form plug in y-2=-9/8(x-2)